Consider the folowing short run production function Q=6L2-0.4L3 how much is labor at each stae of production
1
Expert's answer
2021-10-13T20:43:35-0400

"Q=6L^2-0.4L^3"

Differentiating Q w.r.t. L:

"\frac{\delta Q}{\delta L}=12L-1.2L^2"

Equate "\frac{\delta Q}{\delta L}=0"

"12L-1.2L^2=0"

Solving the equation quadratically, we get:

"L=\frac{-12}{-2.4}"

"L=5."